OpenStudy (anonymous):

2x - y = 5 x + 3y = 7 What is the value of the system determinant? 5 6 7

OpenStudy (anonymous):

x=7-3y according to the bottom equation. Plug that in to the top equation to solve it

OpenStudy (anonymous):

how do i plug it in??

OpenStudy (anonymous):

If the equation is 2(x)-y=5 and x=7-3y then 2(7-3y)-y=5 Does that make sense to you?

OpenStudy (anonymous):

where did the x go

OpenStudy (anonymous):

A drawing might help you one minute

OpenStudy (anonymous):

 The second equation can be rearranged to make x equal to 7-3y. Because now we know x is equal to that, we can replace x in the first equation with the rearranged equation. This is called substitution